# Uncovering the Functional Roles of Nonlinearity in Memory

### v1

@@ Abstract
Memory and long-range temporal processing are core requirements for sequence modeling tasks across natural language processing, time-series forecasting, speech recognition, and control. While nonlinear recurrence has long been viewed as essential for enabling such mechanisms, recent work suggests that linear dynamics may often suffice. In this study, we go beyond performance comparisons to systematically dissect the functional role of nonlinearity in recurrent networks—identifying both when it is computationally necessary, and what mechanisms it enables. We use Almost Linear Recurrent Neural Networks (AL-RNNs), which allow fine-grained control over nonlinearity, as both a flexible modeling tool and a probe into the internal mechanisms of memory. Across a range of classic sequence modeling tasks and a real-world stimulus selection task, we find that minimal nonlinearity is not only sufficient but often optimal, yielding models that are simpler, more robust, and more interpretable than their fully nonlinear or linear counterparts. Our results provide a principled framework for selectively introducing nonlinearity, bridging dynamical systems theory with the functional demands of long-range memory and structured computation in recurrent neural networks, with implications for both artificial and biological neural systems.
@@ 1 Introduction
Modeling sequences with temporal dependencies remains a core challenge in machine learning, vital for domains such as language, speech, control, and time-series forecasting [28, 65, 35, 36]. These tasks often depend on long-range memory—requiring models to store, transform, and retrieve information across time. Classical results have long pointed to the necessity of nonlinear recurrence for such operations, showing that recurrent neural networks (RNNs) with nonlinear dynamics are both universal approximators and Turing complete [59].
@@
Yet, recent work has revived interest in linear recurrent models, particularly due to their computational advantages. Linear systems support parallel computation and scale efficiently [43], and when equipped with expressive input and output mappings, they have been shown to perform surprisingly well on long-horizon tasks [24, 48]. These findings revive a deeper question: To what extent is nonlinearity necessary for temporal computation, and what role does it actually play?
@@
This question is especially relevant when considering interpretability. Linear systems, despite their dynamical limitations [61], are analytically tractable and structurally transparent, making them appealing in settings where understanding the internal dynamics of a model is as important as its raw performance (as, for instance, in neuroscience [62, 16]). Nonlinear systems, by contrast, afford richer dynamics, but typically introduce a higher degree of complexity.
Reflecting this trade-off, the landscape of sequence models has grown into a diverse collection of architectures combining linearity and nonlinearity in different ways [50], for instance through mechanisms such as input-dependent parameter nonlinearities [22] or gating mechanisms [44]. These innovations, while empirically successful, often blur the functional roles of linear and nonlinear components, obscuring their individual contributions to sequence modeling. This lack of clarity motivates our investigation: we seek to disentangle how and when nonlinearity is functionally necessary, and to better understand its computational benefits across diverse tasks.
We tackle this question using the Almost-Linear RNN (AL-RNN), a model initially proposed for dynamical systems reconstruction (DSR) [9]. Here, we repurpose the AL-RNN as a natural tool to investigate a wide class of sequence modeling tasks. The AL-RNN is uniquely suited for this purpose due to its ability to adjust the degree of nonlinearity in its recurrence, allowing for a systematic exploration of where and when nonlinearity is functionally necessary. After inference, the AL-RNN offers dynamical and computational interpretability: First, its piecewise linear (PWL) structure allows analytical access to its dynamical objects such as fixed points, stability properties, and bifurcations [17]. Second, the AL-RNN naturally generates a symbolic representation of its internal state transitions, structured through a bitcode that reflects its nonlinear switching mechanisms [9].
@@
In our study, we apply the AL-RNN across a diverse set of classic memory tasks, including neural recordings from a rodent performing a stimulus selection task, systematically exploring scenarios where linearity suffices, minimal nonlinearity is beneficial, or strong nonlinearity is required. A central finding of our study is that sparsely applied nonlinearity—introduced selectively within an otherwise linear system—can match or sometimes even exceed the performance of fully nonlinear models while retaining strong interpretability. In doing so, we reveal task-specific nonlinear mechanisms such as gating, rule-based integration, and memory-dependent transients, all embedded within a mostly linear dynamical backbone. Our study sheds light on the minimum necessary conditions for memory in recurrent models, offering a principled framework guiding the design of sequence models that are both computationally efficient and analytically tractable.
@@ 2 Related Work
The ability to represent, retain, and manipulate memory over time has been a central challenge in sequence modeling. Early RNNs, such as those proposed by Elman [18] and Jordan [30], introduced the paradigm of learning temporal dependencies through hidden state recurrence. However, these simple RNNs struggled to capture long-range dependencies due to vanishing and exploding gradients [6]. This limitation was addressed by architectures like Long Short-Term Memory (LSTM) networks [28] and Gated Recurrent Units [13]. Such gated RNNs, and more recent extensions such as Long Expressive Memory [56], established gated recurrence as a cornerstone of sequence modeling, supporting stable memory retention across extended time spans. These models successfully mitigated gradient issues, but the introduction of nonlinearity through gating also increased the complexity of their internal state dynamics. To address this, an alternative line of work has focused on combining the expressivity of nonlinear dynamics with the analytical tractability of linear models. This gave rise to piecewise linear and switching models—such as switching linear dynamical systems [19, 38, 37] and piecewise linear recurrent neural networks (PLRNNs) [15, 8, 9]. These architectures segment the state space into locally linear regions, enabling rich dynamical behaviors through structured transitions between subspaces while preserving mathematical tractability.
@@
While the computational complexity of RNNs scales with $\mathcal{O}(T)$, $T$ being the sequence length, the inherently sequential nature of RNNs makes their training inefficient to parallelize on modern GPU hardware. The Transformer architecture [65] has therefore largely replaced recurrence with global self-attention across many sequence-modeling tasks. Self-attention enables parallel sequence processing without relying on a persistent hidden state. This parallelism, however, comes at the cost of computational efficiency, scaling with $\mathcal{O}(T^{2})$. Further, it models memory only within a finite context window. As a result, Transformers lack causal, evolving memory, making them harder to interpret and less inherently efficient for very long-range dependencies [50].
@@
This quadratic scaling has motivated the exploration of more efficient alternatives for long-sequence modeling. Recent interest has focused on structured linear state-space models (SSMs). These models implement linear recurrence, which can be unrolled using spectral methods or convolutions. Advanced parallel scanning techniques, such as Blelloch’s scan [7], enable linear RNNs to achieve highly efficient parallelization, reducing complexity to $\mathcal{O}(\log T)$ [43]. Theoretical advancements have even demonstrated that deep linear RNNs with nonlinear mixing layers are universal approximators of regular sequence maps, broadening their applicability to complex temporal tasks [48, 49]. Modern SSMs like S4 [24], grounded in HiPPO theory [23], leverage these principles to match, or even surpass, Transformers on long-range sequence benchmarks [63]. Recent extensions have introduced more efficient initialization schemes and streamlined designs [25, 60, 26], enhancing both stability and performance. Some models, like Mamba [22] and Gated SSMs [44], reintroduce nonlinearity through structured gating mechanisms while preserving the parallelizable structure of SSMs.
@@ 3 Method
To dissect the roles of linear and nonlinear recurrence, we adopt an SSM-inspired architecture: the recently proposed AL-RNN [9]. In this model, latent dynamics evolve under a combination of linear and PWL transition functions, modulated by external inputs. Nonlinearities are applied only to a subset $P\leq M$ of the hidden state, making AL-RNNs a simplified variant of piecewise linear RNNs [15, 8].
An AL-RNN is defined by:
@@
with $\Phi^{*}(\bm{z}_{t})=\left[z_{1,t},\cdots,z_{M-P,t},\max(0,z_{M-P+1,t}),\cdots,\max(0,z_{M,t})\right]^{T}$, and $\bm{z}_{t}\in\mathbb{R}^{M}$ being the system’s latent state. To avoid redundant parameterization, we define $\bm{A}\in\mathbb{R}^{M\times M}$ as a diagonal matrix where entries corresponding to the linear (first $M-P$) units are set to zero, $\bm{A}=\mathrm{diag}(0,\ldots,0,a_{M-P+1},\ldots,a_{M}),$ so that linear self-connections are only assigned to the nonlinear units. The matrix $\bm{W}\in\mathbb{R}^{M\times M}$ captures interactions based on the partially nonlinear transformed state $\Phi^{*}(\bm{z})$, $\bm{C}\in\mathbb{R}^{M\times K}$ weighs the K-dimensional external inputs $\bm{s}_{t}$, and $\bm{h}\in\mathbb{R}^{M}$ is a bias term.
@@
The AL-RNN partitions the state space into $2^{P}$ linear subregions separated by switching boundaries, within which the dynamics are linear and analytically tractable (see Fig. 9). Each linear subregion of the AL-RNN can be uniquely identified by a binary bitcode of length $P$, corresponding to the on/off (positive/negative) activation state of the $P$ ReLU units. Since the ReLU nonlinearity partitions each dimension at zero, this yields $2^{P}$ possible configurations of active units. By tracking which bitcodes occur during inference, we can quantify and visualize how many of these subregions are actually used by the network, offering a compact representation of its functional complexity.
This PWL structure offers a key advantage: it preserves linear dynamics within each subregion while still enabling the network to express nonlinear behavior through discrete transitions between linear subregions. This allows for efficient computation and clear interpretability. Transitions between subregions are explicitly detectable, and within each regime, the system’s dynamics can be analytically characterized in terms of fixed points and stability properties [15, 17] (see Appx. A.2). To retain interpretability, we constrain our models to a single AL-RNN layer throughout.
@@
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we incorporate a regularization term based on Schmidt et al. [57] (see Appx. A.1.2 for details and performance comparisons). Across tasks, the AL-RNN is embedded in an encoder–decoder framework: raw sequences $\{\bm{x}_{t}\}_{t=1:T}$ are first mapped to latent inputs $\{\bm{s}_{t}\}_{t=1:T}$ via task-specific encoders. When the input modality has known structure (e.g., audio, vision, language), we apply appropriate nonlinear preprocessing (e.g., spectral representations, convolutions, or word embeddings, see Appx. A.1.3). For symbolic inputs, we directly use one-hot encodings. The AL-RNN processes these inputs, and task outputs are decoded from the latent states $\{\bm{z}_{t}\}_{t=1:T}$ or the final state $\bm{z}_{T}$ via a linear readout layer. While sophisticated nonlinear decoders could potentially offload complexity from the recurrent model, this consistent use of linear decoding isolates the contribution of recurrent dynamics to memory and task performance, ensuring that any long-range dependencies and nonlinear effects are captured within the recurrent structure itself.
Memory Mechanisms in RNNs
Sequential tasks require recurrent systems to store and manipulate information over time. RNNs implement these memory functions through distinct dynamical mechanisms that govern the evolution of hidden states, and associated computational mechanisms that determine how information is processed.
Linear RNNs exhibit a restricted but analytically tractable set of dynamics, governed by the spectral properties of the recurrence matrix [61]. Slow modes (eigenvalues near the unit circle) allow information to persist over time, while marginal stability supports long-lived activity in finely tuned cases. Oscillatory dynamics, from complex eigenpairs, enable rhythmic trajectories but require precise spectral alignment. The system is time-invariant and input-linear, lacking state-dependent transitions. These dynamics support basic memory operations such as leaky integration, graded retention, and temporal filtering [58, 54, 21]. Oscillatory modes - although fragile - can encode phase-sensitive information [27]. Linear RNNs cannot implement gating or context-sensitive routing, but can approximate input projection and superposition, enabling limited compositional processing.
Nonlinear RNNs support a broad repertoire of dynamical regimes, including multistability, where multiple attractors allow for distinct, stable internal states. Thresholded activations act as switches, enabling state-dependent transitions. More complex dynamics such as chaotic itinerancy and metastability generate structured, non-convergent trajectories. These dynamics support associative memory [29, 4], task and context-dependent switching [5, 67], and context-dependent routing, where identical inputs lead to different outcomes depending on internal state. Nonlinear gating allows selective integration or suppression of information [2]. Transient trajectories support sequential processing and structured recall [52, 64].

### v2

# Uncovering the Computational Roles of Nonlinearity in Sequence Modeling Using Almost-Linear RNNs

@@ Abstract
Sequence modeling tasks across domains such as natural language processing, time-series forecasting, speech recognition, and control require learning complex mappings from input to output sequences. In recurrent networks, nonlinear recurrence is theoretically required to universally approximate such sequence-to-sequence functions; yet in practice, linear recurrent models have often proven surprisingly effective. This raises the question of when nonlinearity is truly required. In this study, we present a framework to systematically dissect the functional role of nonlinearity in recurrent networks— allowing to identify both when it is computationally necessary, and what mechanisms it enables. We address the question using Almost Linear Recurrent Neural Networks (AL-RNNs), which allow the recurrence nonlinearity to be gradually attenuated and decompose network dynamics into analyzable linear regimes, making the underlying computational mechanisms explicit. We illustrate the framework across a diverse set of synthetic and real-world tasks, including classic sequence modeling benchmarks, an empirical neuroscientific stimulus-selection task, and a multi-task suite. We demonstrate how the AL-RNN’s piecewise linear structure enables direct identification of computational primitives such as gating, rule-based integration, and memory-dependent transients, revealing that these operations emerge within predominantly linear dynamical backbones. Across tasks, sparse nonlinearity plays several functional roles: it improves interpretability by reducing and localizing nonlinear computations, promotes shared (rather than highly distributed) representations in multi-task settings, and reduces computational cost by limiting nonlinear operations. Moreover, sparse nonlinearity acts as a useful inductive bias: in low-data regimes, or when tasks require discrete switching between linear regimes, sparsely nonlinear models often match or exceed the performance of fully nonlinear architectures. Our findings provide a principled approach for identifying where nonlinearity is functionally necessary in sequence models, guiding the design of recurrent architectures that balance performance, efficiency, and mechanistic interpretability.
@@ 1 Introduction
Modeling sequences with temporal dependencies remains a core challenge in machine learning, with applications spanning natural language processing, speech recognition, video analysis, control, and forecasting (Hochreiter and Schmidhuber, 1997, Vaswani et al., 2017, Lillicrap et al., 2019, Lim et al., 2021). Sequence modeling tasks often depend on long-range memory (Bengio et al., 1994), requiring models to store, transform, and retrieve information across time. Classical results have long pointed to the necessity of nonlinear recurrence for such operations, showing that recurrent neural networks (RNNs) with nonlinear dynamics are both universal approximators and Turing complete (Siegelmann and Sontag, 1995).
@@
Yet, recent work has revived interest in linear recurrent models, particularly due to their computational advantages (Orvieto et al., 2023). Linear systems support parallel computation and scale efficiently (Martin and Cundy, 2018), and - when equipped with expressive input and output mappings - perform surprisingly well on long-horizon tasks (Gu et al., 2022, Orvieto et al., 2023). These findings revive a deeper question: How can we systematically assess to what extent nonlinearity is necessary for temporal computation, and what role it plays?
@@
This question is especially relevant when considering interpretability. Linear systems, despite their dynamical limitations (Strogatz, 2018), are analytically tractable and structurally transparent, making them appealing in settings where understanding the internal dynamics of a model is as important as its raw performance (as, for instance, in neuroscience; Sussillo and Barak (2013), Durstewitz et al. (2023)). Nonlinear systems, by contrast, afford richer dynamics, but typically introduce a higher degree of complexity. Reflecting this trade-off, the landscape of sequence models has grown into a diverse collection of architectures combining linearity and nonlinearity in different ways (Patro and Agneeswaran, 2024), for instance through mechanisms such as input-dependent parameter nonlinearities (Gu and Dao, 2023) or gating mechanisms (Mehta et al., 2023). These innovations, while empirically successful, often blur the functional roles of linear and nonlinear components, obscuring their individual contributions to sequence modeling.
Motivated by this, we aim to provide a framework that disentangles how, when, and if nonlinearity is functionally necessary, and to demonstrate its utility for uncovering the computational mechanisms underlying sequence modeling. We tackle this question using the Almost-Linear RNN (AL-RNN), a model initially proposed for dynamical systems reconstruction (Brenner et al., 2024a). Here, we repurpose the AL-RNN as a natural tool to investigate a wide class of sequence modeling tasks. It is uniquely suited for this purpose because it makes nonlinearity both controllable and observable. Built from both linear and piecewise-linear units, the AL-RNN allows to tune the strength of nonlinear recurrence by varying the number of nonlinear units. At the same time, its piecewise-linear structure partitions the hidden state space into distinct linear subregions. Tracking which subregion is active over time yields a symbolic bitcode trace of the computation, which we use to analyze the switching structure that implements temporal processing. Finally, the linear dynamics within each subregion enable complementary dynamical analyses (e.g., fixed points and stability) (Eisenmann et al., 2024), connecting the bitcode-level view to interpretable dynamical mechanisms.
@@
We apply the AL-RNN across a diverse set of classic memory benchmarks, a multi-task setting (Driscoll et al., 2024), as well as an empirical dataset of neural recordings from a rodent performing a stimulus selection task, using each setting to demonstrate how nonlinear mechanisms can be systematically identified, while exploring scenarios where linearity suffices, minimal nonlinearity is beneficial, or strong nonlinearity is required. In doing so, we reveal task-specific nonlinear mechanisms such as gating, rule-based integration, and memory-dependent transients, often embedded within a mostly linear dynamical backbone.
In most single task settings, reducing nonlinearity matches or even exceeds the performance of fully nonlinear models while retaining interpretability and making computational motifs and mechanisms directly traceable. In the multi-task setting, it forces reuse of a small set of nonlinear computational motifs, promoting shared representations across tasks, improving sample efficiency and exposing a mechanistic map of cross-task structure. Collectively, our results suggest that sparse nonlinearity provides a useful inductive bias that enables the emergence of nonlinear motifs where needed, while preserving predominantly linear dynamics. Overall, our study sheds light on the minimum necessary conditions for memory in AL-RNNs and related architectures, offering a principled framework guiding the design of sequence models that are both computationally efficient and analytically tractable.
@@ 2 Related Work
The ability to represent, retain, and manipulate memory over time has been a central challenge in sequence modeling. Early RNNs, such as those proposed byElman (1990) and Jordan (1997), introduced the paradigm of learning temporal dependencies through hidden state recurrence. However, these simple RNNs struggled to capture long-range dependencies due to vanishing and exploding gradients (Bengio et al., 1994). This limitation was addressed by architectures like Long Short-Term Memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Units (Chung et al., 2014). Such gated RNNs, and more recent extensions such as Long Expressive Memory (Rusch et al., 2022), established gated recurrence as a cornerstone of sequence modeling, supporting stable memory retention across extended time spans. These models successfully mitigated gradient issues, but the introduction of nonlinearity through gating also increased the complexity of their internal state dynamics. To address this, an alternative line of work has focused on combining the expressivity of nonlinear dynamics with the analytical tractability of linear models. This gave rise to piecewise linear and switching models—such as switching linear dynamical systems (Fox et al., 2008, Linderman et al., 2016; 2017) and piecewise linear recurrent neural networks (PLRNNs) (Durstewitz, 2017, Brenner et al., 2022; 2024a). These architectures segment the state space into locally linear regions, enabling rich dynamical behaviors through structured transitions between subspaces while preserving mathematical tractability.
@@
While the computational complexity of RNNs scales with $\mathcal{O}(T)$, $T$ being the sequence length, the inherently sequential nature of RNNs makes their training inefficient to parallelize on modern GPU hardware. The Transformer architecture (Vaswani et al., 2017) has therefore largely replaced recurrence with global self-attention across many sequence-modeling tasks. Self-attention enables parallel sequence processing without relying on a persistent hidden state. This parallelism, however, comes at the cost of computational efficiency, scaling with $\mathcal{O}(T^{2})$. Further, it models memory only within a finite context window. As a result, Transformers lack causal, evolving memory, making them harder to interpret and less inherently efficient for very long-range dependencies (Patro and Agneeswaran, 2024).
@@
This quadratic scaling has motivated the exploration of more efficient alternatives for long-sequence modeling. Recent interest has focused on structured linear state-space models (SSMs). These models implement linear recurrence, which can be unrolled using spectral methods or convolutions. Advanced parallel scanning techniques, such as Blelloch’s scan (Blelloch, 1990), enable linear RNNs to achieve highly efficient parallelization, reducing complexity to $\mathcal{O}(\log T)$ (Martin and Cundy, 2018). Theoretical advancements have even demonstrated that deep linear RNNs with nonlinear mixing layers are universal approximators of regular sequence maps, broadening their applicability to complex temporal tasks (Orvieto et al., 2023; 2024). Modern SSMs like S4 (Gu et al., 2022), grounded in HiPPO theory (Gu et al., 2020), leverage these principles to match, or even surpass, Transformers on long-range sequence benchmarks (Tay et al., 2021). The Legendre Memory Unit (LMU) (Voelker et al., 2019) similarly enforces a fixed linear memory subspace derived from continuous-time dynamics, combined with a nonlinear update. Recent extensions have introduced more efficient initialization schemes and streamlined designs (Gupta et al., 2022, Smith et al., 2023, Hasani et al., 2023), enhancing both stability and performance. Some models, like Mamba (Gu and Dao, 2023) and Gated SSMs (Mehta et al., 2023), reintroduce nonlinearity through structured gating mechanisms while preserving the parallelizable structure of linear SSMs. Sequence models also increasingly incorporate hybrid architectures that blend recurrent, convolutional, and attention mechanisms (Peng, 2023, Qin et al., 2023, Poli et al., 2023, Lieber et al., 2024, Sieber et al., 2024). However, this proliferation of architectures, driven primarily by performance and efficiency, makes mechanistic interpretation increasingly challenging, as the structural assumptions each model encodes often remain implicit.
@@ 3 Method
To dissect the roles of linear and nonlinear recurrence, we adopt an SSM-inspired architecture: the recently proposed AL-RNN (Brenner et al., 2024a). In this model, latent dynamics evolve under a combination of linear and PWL transition functions, modulated by external inputs. Nonlinearities are applied only to a subset $P\leq M$ of the hidden state, making AL-RNNs a simplified variant of piecewise linear RNNs (Durstewitz, 2017, Brenner et al., 2022). An AL-RNN is defined by:
@@
where $\bm{z}_{t}\in\mathbb{R}^{M}$ is the system’s latent state. The function $\Phi^{*}(\cdot)$ applies a nonlinearity selectively to only the last $P$ units:
$\Phi^{*}(\bm{z}_{t})=\left[z_{1,t},\cdots,z_{M-P,t},\phi(z_{M-P+1,t}),\cdots,\phi(z_{M,t})\right]^{T},$ (2)
@@
where $\phi(\cdot)$ is a scalar nonlinearity. The original AL-RNN formulation (Brenner et al., 2024a) used the ReLU nonlinearity $\phi(z)=\max(0,z)$. While we also tested other nonlinearities including GeLU, tanh, and hardtanh, only PWL functions like ReLU preserve the property that the system operates as a composition of linear dynamical systems. To avoid redundant parameterization, we define $\bm{A}\in\mathbb{R}^{M\times M}$ as a diagonal matrix where entries corresponding to the linear (first MPM-P) units are set to zero, $\bm{A}=\mathrm{diag}(0,\ldots,0,a_{M-P+1},\ldots,a_{M}),$ so that linear self-connections are only assigned to the nonlinear units. The matrix $\bm{W}\in\mathbb{R}^{M\times M}$ captures interactions based on the partially nonlinear transformed state $\Phi^{*}(\bm{z})$, $\bm{C}\in\mathbb{R}^{M\times K}$ weighs the K-dimensional external inputs $\bm{s}_{t}$, and $\bm{h}\in\mathbb{R}^{M}$ is a bias term.
Figure 1: Illustration of the AL-RNN and bitcode assignment. The example displays a 6-dimensional AL-RNN with four linear and two PWL units (left). The PWL units partition the state space into four linear subregions (right). Bitcodes encode positive $(1)$ and negative $(0)$ activation values for these units. Each subregion corresponds to a unique bitcode ($0$, 01, 10, 11) and is governed by a distinct linear dynamical system (DS) with its own recurrence matrix $\mathbf{W}_{\Omega^{i}}$. An example trajectory (bottom) traverses a sequence of these subregions over time, experiencing discrete switches in dynamics when crossing subregion boundaries (marked by dashed vertical lines). Within each subregion, the dynamics remain linear.
Computational Mechanisms in RNNs
Sequential tasks require recurrent networks to store information over time and to manipulate it in ways that depend on both past states and incoming inputs. These functions can be understood through the dynamics that govern hidden states and the computations they enable, implementing objects such as slow modes, attractors, and transient trajectories. From a computational perspective, these structures implement operations such as accumulation, recall, gating, and context-dependent routing.
Linear RNNs exhibit a restricted but analytically tractable set of dynamics, governed by the spectral properties of the recurrence matrix (Strogatz, 2018). Autonomous dynamics are limited to fixed points and modal decompositions: slow modes (eigenvalues near the unit circle), which support information storage through leaky integration or graded retention, complex eigenpairs, which generate oscillatory modes for rhythmic or phase-sensitive coding, and unstable modes, which can produce exponential growth or decay depending on gain and initialization (Seung, 1996, Ratcliff and McKoon, 2008, Gold and Shadlen, 2007, Henaff et al., 2016). These autonomous motifs can support stable (or deliberately unstable) memory traces but, because the update is globally linear, do not by themselves implement state-dependent reconfiguration. Input-driven dynamics in linear RNNs therefore reduce to uniform accumulation or filtering, lacking mechanisms for gating, selective recall, or context-dependent switching. A key strength lies in their analytic transparency and stability when dominated by slow modes, but their computational repertoire is fundamentally limited compared to nonlinear models.
Nonlinear RNNs differ both in terms of the autonomous dynamics they can generate, and how they can integrate external inputs. Autonomous nonlinear recurrence enables multistability, where multiple attractors allow for distinct, stable internal states that support e.g. associative memory (Hopfield, 1982, Amari, 1977). More complex motifs such as k-cycles, chaotic attractors, itinerancy, and metastability generate structured trajectories that do not converge to a single fixed point, supporting sequential processing and structured recall (Rabinovich et al., 2008, Tsuda, 2015). In terms of input integration, thresholded activations can act as switches, enabling state-dependent transitions and context-dependent routing (Balaguer-Ballester et al., 2011, Wang, 2008). This allows identical inputs to produce different outcomes depending on the internal state. Nonlinear gating, in turn, selectively integrates or suppresses information (Ackley et al., 1985). Nonlinear RNNs therefore excel when tasks require flexible memory or complex temporal structure, but their richer dynamics also make them harder to infer, analyze, and track compared to linear models.
@@
Piecewise Linear RNNs aim to reconcile these trade-offs by combining the analytical tractability of linear dynamics with the computational expressivity of nonlinear models. From a dynamical systems perspective, piecewise linear systems offer a principled and general framework: any sufficiently smooth nonlinear DS can be approximated arbitrarily well by partitioning the state space into linear subregions (Storace and De Feo, 2004). The AL-RNN, which we employ here, implements the simplest version of this idea that can be learned automatically from data in an end-to-end fashion. It makes no assumptions about the switching mechanism itself, and requires no prior specification of how many subregions are needed or where transitions should occur. The AL-RNN partitions the state space into $2^{P}$ linear subregions separated by switching boundaries, within which the dynamics are linear and analytically tractable (see Fig. 1). Each linear subregion of the AL-RNN can be uniquely identified by a binary bitcode of length $P$, corresponding to the on/off (positive/negative) activation state of the $P$ ReLU units (Fig. 1,right). Since the ReLU nonlinearity partitions each dimension at zero, this yields $2^{P}$ possible configurations of active units. By tracking which bitcodes occur during inference, we can quantify and visualize how many of these subregions are actually used by the network, offering a compact representation of its functional complexity. This PWL structure thus inherits the advantages from both paradigms: the analytic transparency of linear systems within each subregion, combined with the representational flexibility of nonlinear models. Further, transitions between subregions are explicitly detectable, and within each regime, the system’s dynamics can be analytically characterized in terms of fixed points and stability properties (Durstewitz, 2017, Eisenmann et al., 2024) (see Appx. A.2).
@@
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we incorporate a manifold regularization term based on Schmidt et al. (2021) (see Appx. A.1.2 for details and performance comparisons). The AL-RNN processes latent inputs $\{\bm{s}_{t}\}_{t=1:T}$ derived from raw inputs $\{\bm{x}_{t}\}_{t=1:T}$ (for details, see Appx. A.1.3), and task outputs are decoded from the latent states $\{\bm{z}_{t}\}_{t=1:T}$ or the final state $\bm{z}_{T}$ via a linear readout layer. This consistent use of linear decoding is central to our interpretability framework. By constraining the decoder to be linear, we ensure that the recurrent dynamics themselves must carry the full computational burden of the task solution. Since the AL-RNN implements a first-order Markov process, this task solution naturally takes the form of a DS, mirroring the classical computational neuroscience perspective on how stateful recurrence encodes computation through evolving trajectories (Durstewitz et al., 2023).